The average age for licensed drivers in a county is µ= 42.6, σ= 12 and the distribution is approximately normal. A county police officer was interested in whether the average age of those receiving parking tickets in the county differed from that of the average age of the population of licensed drivers in the county. She obtained a random sample of N = 25 drivers in the county, who received parking tickets. The average age for these drivers was M= 40.5.
1. Identify the appropriate hypothesis for analyzing this data including the conditions for using this test.
2. Complete the four steps of hypothesis testing necessary to determine whether this group differs from the population of drivers in the county and include each step in your final study group answer.
3. Given the conclusion found in #2, which type of error (Type I or Type II) is a possibility? Why?
4. Given the possible type of error (Type I or Type II) made in this situation, what could researchers do to reduce the risk of this error?
Thank you!

Answer by stanbon(75887)

The average age for licensed drivers in a county is µ= 42.6, σ= 12 and the distribution is approximately normal. A county police officer was interested in whether the average age of those receiving parking tickets in the county differed from that of the average age of the population of licensed drivers in the county. She obtained a random sample of N = 25 drivers in the county, who received parking tickets. The average age for these drivers was M= 40.5.
-----------
Ho: ua = 42.6
Ha: ua # 42.6 (claim)
----
x-bar = 40.5
t(40.5) = (40.5-42.6)/[12/sqrt(25)] = -2.1*5/12 = -0.875
p-value = 2*P(t<-0.875 when df = 24) = 0.39
----
Conclusion:: Since the p-value is greater that 5%, fail to reject
Ho at the 5% confidence level.
Conclusion: The test results do not support the claim.
----------------------------
1. Identify the appropriate hypothesis for analyzing this data including the conditions for using this test.
2. Complete the four steps of hypothesis testing necessary to determine whether this group differs from the population of drivers in the county and include each step in your final study group answer.
3. Given the conclusion found in #2, which type of error (Type I or Type II) is a possibility? Why?
4. Given the possible type of error (Type I or Type II) made in this situation, what could researchers do to reduce the risk of this error?
-------
Cheers,
Stan H.
---------------